Runtan Inc. has just paid an annual dividend of $0.45 per share. Analysts expect the firm's dividends to grow by 5% forever. Its stock price is $37.4 and its beta is 1.3. The risk-free rate is 2% and the expected return on the market portfolio is 8%.

Part 1

What is the best guess for the cost of equity?
Answer #1

Ke = [ D1 / P0 ] + g

D1 = D0 (1+g)

= 0.45 ( 1 + 0.05 )

= 0.45 ( 1.05 )

= $ 0.4725

D1 = Expected Div after 1 Year

P0 = Current price

g = Growth Rate

Ke = [ D1 / P0 ] + g

= [ 0.4725 / 37.4 ] + 0.05

= 0.0126 + 0.05

= 0.0626 i.e 6.26%

Best guess of cost of Equity is 6.26%